a calorimeter contained 75.0g of water at 16.95 degress celcius. a 93.3 g sample of iron at 65.58 degrees celcius was placed in it, giving a final temperature of 19.68 degres celius for the system. calculate the heat capacity of the calorimeter specific heats are 4.184J/g/degrees celcius for h20 and 0.444 J?g?degrees celicius for Fe
1
Expert's answer
2019-06-21T04:16:55-0400

"\Delta t=" 65.58 - 19.68 = 45.9oC

"q=mc\Delta t =93.3g*45.9C*0.444 J/g*C=1901.42 J"

"\Delta t(H2O)= 19.68 - 16.95 = 2.73"oC

"q=75g*4.184J\/g*C*2.73C=856.674 J"

1901.42 J - 856.674 J = 1044.746 J

"c=q/ \Delta t =1044.746 J/2.73"oC=382.69 J/oC